OpenStudy (anonymous):

Divide. 48divided by 15.6 using long divison

myininaya (myininaya):

so 48/15.6=480/156 if we multiply the top and bottom by 10

myininaya (myininaya):

how many times can 156 go into 480

myininaya (myininaya):

156*4 is too much so 156*3 we do just fine

myininaya (myininaya):

156*3=468

myininaya (myininaya):

do 480-468

OpenStudy (anonymous):

12

myininaya (myininaya):

now bring down a 0 from 480.0

myininaya (myininaya):

and ask yourself how many times does 156 go into 120

OpenStudy (anonymous):

1

myininaya (myininaya):

no its 0

myininaya (myininaya):

so you bring down another 0

myininaya (myininaya):

after putting that 0 where the quotient is

myininaya (myininaya):

how far are we suppose to go as in the length of the decimal number

myininaya (myininaya):

this number is forever

myininaya (myininaya):

do we stop at the hundredths spot

OpenStudy (anonymous):

no

myininaya (myininaya):

we need a stopping spot because like i said this number has no end

OpenStudy (anonymous):

3.0769230769230 after that it just repeating itself over and over again

myininaya (myininaya):

oh yeah it does have a pattern

myininaya (myininaya):

you just put a bar over the repeating part

OpenStudy (anonymous):

3.0769230769230769230769230 .....on and on forever :) you'd never get to the end of that :)... and yes to the bar.. good job :)